# Transfer Learning for Spatial Autoregressive Models

### v1

@@ 6 Conclusion
We proposed a new transfer learning method under the SAR models, aiming to address the challenges of small sample sizes and spatial dependencies. We derived the theoretical convergence rates for a two-step $\mathcal{A}$-TranSAR estimator when the informative source data sets are known, which is more efficient than the traditional SAR estimator as long as the sample size of the informative source data sets is large enough. Then, we introduced an algorithm for detecting the informative source datasets based on spatial residual bootstrap and proved its detection consistency. The empirical application of this methodology in the prediction of U.S. election results in swing states utilizing demographic and geographical data demonstrates its potential to improve the traditional spatial methodology. Further research might focus on tests for the detection of informative sets and improving the accuracy of coefficients estimation within the transfer learning framework.
Acknowledgements
@@
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT 4.0 in order to polish the writing. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.

### v2

@@ 7 Conclusion
To more accurately explore the spatial factors influencing swing states, we propose a method to enhance the estimation of spatial model in swing states by transferring knowledge from other states. We introduce a transfer learning method within the framework of SAR models, aimed at addressing spatially dependent data of small sample size. The empirical application of this methodology, specifically in predicting U.S. election results in swing states using demographic and geographical data, demonstrates its potential to enhance traditional spatial methods. Based on our comprehensive analysis and the application of the tranSAR model, our prediction for the 2024 U.S. presidential election is a victory for the Democratic party. Future research could focus on developing tests to detect informative sets and improving the accuracy of coefficient estimation within the transfer learning framework.
Supplementary Materials
All technical proofs of theorems and the additional simulations are included in the online supplementary materials.
Funding